Find the partial fraction decomposition for each of the following rational expressions. Then use the partial fraction to give an antiderivative for each function.

(2-6) (3 + 2)2
Answer #1

sot: x (2-6) (x+2) 2 using pavetial fraction method (2-6) (x+2) 2 -07 2242) + (242) X A (x+2)² + B(x-6) (x-6) (x+2)² o l +2)(x+6) (+2) - = (A+C) x2 + (4A+B-40) x + (4A-6B-126) (x-c) (X+2) 2 en compairing both the sides, we get, » A+C = 0 ->0. → 4A+B→ 640 = -6 C = -6 64 C = - 3 3211 i A= -6 And. B = 1+80 = 1+8 [3] * I+ (z] in Replacing each of the coeffivents partial fractNow, entegerating wert a weget, Ja-6)(x+2) (x-6) (x+3) a (x+2) de da= lnxto 10(+6) ++/6+2)* dan 10 (4+2) + c [: J4d2 = tnx +c